OpenStudy (anonymous):

Simplify (6ab3c)(-abc2).

OpenStudy (anonymous):

Just multiply them..

OpenStudy (zeesbrat3):

\[-6a ^{2}b ^{4}c ^{3} \] if I'm not mistaken

OpenStudy (anonymous):

\[(6ab^3c)(-abc^2)\]

OpenStudy (anonymous):

Yup. \[(6ab^{3}c)(-abc^{2}) = -6a^{1 + 1}\space b^{3 + 1}\space c^{1 + 2} = ?\]

OpenStudy (anonymous):

what about \[\sqrt{_{18}^{7}}\]

OpenStudy (anonymous):

Did you mean? \[\sqrt[7]{18}\] ??

OpenStudy (anonymous):

https://media.glynlyon.com/g_alg01_2011/8/test8.gif

OpenStudy (anonymous):

See: \[\sqrt{18} = 3\sqrt{2}\] \[\sqrt{\frac{a}{b}} = \frac{\sqrt{a}}{\sqrt{b}}\]

OpenStudy (anonymous):

sooo its https://media.glynlyon.com/g_alg01_2011/8/sr6_14.gif